OpenStudy (idealist10):

Solve the initial-value problem y'=2y-y^2, y(0)=1.

OpenStudy (idealist10):

Here's the work: dy/dx=2y-y^2 dy/(2y-y^2)=dx A/y+B/(2-y)=1 A=1/2, B=1/2, (1/2)ln abs(y)+(1/2)ln abs(2-y)=x+c sqrt(y)*sqrt(2-y)=x+c sqrt(2y-y^2)=x+c 2y-y^2=(x+c)^2 -y^2+2y=(x+c)^2 -(y^2-2y)=(x+c)^2 (y^2-2y+1-1)=-(x+c)^2 (y-1)^2=1-(x+c)^2 y-1=+/-sqrt(1-(x+c)^2) y=1+/-sqrt(1-(x+c)^2) y(0)=1+/-sqrt(1-c^2)=1

OpenStudy (idealist10):

The answer should be y=2/(1+e^(-2x)).

OpenStudy (anonymous):

\[\int\frac{dy}{2-y}={-}\ln|2-y|+C\]

OpenStudy (anonymous):

So you should have \[\frac{1}{2}\ln|y|-\frac{1}{2}\ln|2-y|=\ln\sqrt{\frac{y}{2-y}}=x+C\] which gives the general solution \[\sqrt{\frac{y}{2-y}}=Ce^x~~\iff~~\frac{y}{2-y}=Ce^{2x}\] Solving for \(y\): \[\begin{align*} y&=2Ce^{2x}-Cye^{2x}\\\\ y(1-Ce^{2x})&=2Ce^{2x}\\\\ y&=\frac{2Ce^{2x}}{1-Ce^{2x}} \end{align*}\]

OpenStudy (idealist10):

Let me check my work.

OpenStudy (idealist10):

I see why. I got (1/2)ln abs(y)+(1/2)ln abs(2-y) but you got - instead of +, is it because of the - in 2-y?

OpenStudy (anonymous):

Yes, you would have substituted \(u=2-y\), which would give \(-du=dy\).

OpenStudy (idealist10):

So I got c=1/3.

OpenStudy (anonymous):

Hold on, I think I made a mistake with the general solution...

OpenStudy (idealist10):

I got y=2e^(2x)/(3-e^(2x)).

OpenStudy (anonymous):

The \[\begin{align*}\\ \frac{y}{2-y}&=Ce^{2x}\\\\ y&=Ce^{2x}-Cye^{2x}&\text{the 2 gets absorbed into the }C\\\\ y(1-Ce^{2x})&=Ce^{2x}\\\\ y&=\frac{Ce^{2x}}{1-Ce^{2x}} \end{align*}\] And so this gives \[\begin{align*} 1&=\frac{C}{1-C}\\ 1-C&=C\\ C&=\frac{1}{2} \end{align*}\] Hmm, WA says this answer is wrong. Let's start from the beginning...

OpenStudy (anonymous):

\[\begin{align*}\frac{dy}{dx}&=2y-y^2\\ \frac{dy}{y(2-y)}&=dx\\ \int\frac{dy}{2y}+\int\frac{dy}{2(2-y)}&=\int dx\\ \frac{1}{2}\ln|y|-\frac{1}{2}\ln|2-y|&=x+C\\ \ln\left|\frac{y}{2-y}\right|&=2x+C\\ \frac{y}{2-y}&=Ce^{2x}\\ y&=2Ce^{2x}-yCe^{2x}\\ y+yCe^{2x}&=2Ce^{2x}\\ y(1+Ce^{2x})&=2Ce^{2x}\\ y&=\frac{2Ce^{2x}}{Ce^{2x}+1} \end{align*}\] I guess the 2 doesn't get absorbed into the \(C\)... Now, \[1=\frac{2C}{C+1}~~\iff~~C=1\] So the solution is \[\large y=\frac{2e^{2x}}{e^{2x}+1}\]

OpenStudy (idealist10):

Finally getting the right answer! Wow, tough problem...Thanks a lot.

OpenStudy (anonymous):

yw